hhanan:

I got 5 or 25 but that is not an option

hhanan:

1 attachment
jhonyy9:

4x^2 -100 = 0 4x^2 = 100 divide both sides by 4 what will get ?

hhanan:

25

jhonyy9:

or using the formula of difference of squares a^2 -b^2 = (a-b)(a+b)

jhonyy9:

yes 25 x^2 = 25 x_1,2 = +/- sqrt25 = ?

hhanan:

5

hhanan:

-25

jhonyy9:

+/- 5

jhonyy9:

bc. 25 = 5*5 or (-5)(-5) yes ?

hhanan:

yes

hhanan:

but the only options are -25 and 25

hhanan:

options: x = 5 or -5 x = 5 x = -5 x = 25 or -25

jhonyy9:

4x^2 -100 = 0 4x^2 = 100 x^2 = 25 x_1,2 = +/- sqrt25 x_1,2 = +/- 5

jhonyy9:

ok ?

jhonyy9:

@Laylalyssa

hhanan:

ok thank you

CocoCammron99:

Oh ok

jhonyy9:

np yw